# Fast and accurate parameter estimation of high-redshift sources with the Einstein Telescope

### v1

@@ II.2 Priors
$\cos$ dec rad $\mathcal{U}(-\pi/2,\pi/2)$
$\sin\theta_{\mathrm{jn}}$ rad $\mathcal{U}(0,\pi)$
@@
The prior distributions $\pi(\theta)$ adopted in Eq. 1 are reported in Tab. 1. In particular, for the detector-frame chirp mass and mass ratio, we choose a prior uniform in primary and secondary masses (Dupletsa, 2025), therefore:
@@
where $d_{\mathrm{L}}\in[5\times 10^{3},~{}5\times 10^{4}]$ Mpc. For the aligned spin component $\chi_{i}$, we form a joint prior as follows:
@@
where $\pi(a_{i})=\mathcal{U}(0,0.9)$ and $\cos\theta_{i}=\mathcal{U}(-1,1)$ (Lange et al., 2018; Ashton et al., 2019).
By convention, we set the Earth orientation and the positions and orientations of the three interferometers to match those at the reference time $t_{\mathrm{ref}}=1126259462.391$ s, which corresponds to the GPS trigger time of GW150914 (Romero-Shaw et al., 2020). The merger time ($t_{\mathrm{geocent}}$) of all events is sampled randomly around this reference time. Fixing $t_{\mathrm{ref}}$ for such short signals is not a loss of generality, since any difference in time can be recovered.
@@ II.3 Bilby
We compare the results obtained with NPE against Bilby (Romero-Shaw et al., 2020), which is a Python library specifically designed for Bayesian inference. Assuming the noise is stationary and Gaussian, the likelihood in Eq. 1 follows (Finn, 1992):
$\mathcal{L}(d|\theta)\propto\exp\left(-\frac{1}{2}(d-h(f,\theta)|d-h(f,\theta))\right),$ (9)
where $h(f,\theta)$ represents the gravitational wave signal as a function of frequency $f$ for a given set of parameters $\theta$, projected onto the antenna pattern of the detector (see Sec. IV.1 for details). Eq. 9 involves the noise-weighted inner product:
$(a|b)=4~{}\mathrm{Re}~{}\int_{f_{\mathrm{min}}}^{f_{\mathrm{max}}}\mathrm{d}f\frac{a^{*}(\theta,f)b(\theta,f)}{S(f)^{2}},$ (10)
where $\mathrm{Re}$ denotes real part and complex conjugate. Moreover, $S(f)$ is the noise ASD, as described in Sec. II.1, which takes into account that GW detectors are not equally sensitive at all frequencies.
To sample from Eq. 1 with the likelihood defined in Eq. 9, we use the nested sampling algorithm nessai (Williams, 2021; Williams et al., 2021, 2023), which offers fast convergence by being accelerated with normalizing flows.
Nested sampling methods produce posterior samples as a result of calculating the evidence $\mathcal{Z}(d)$ (Skilling, 2004, 2006; Sivia and Skilling, 2006; Buchner, 2021; Ashton et al., 2022). Initially, a set of live points is drawn from the prior distribution. During each iteration, the live point with the lowest likelihood is replaced by a new sample. The algorithm ends once the remaining prior volume contributes less than a predefined fraction of the evidence (e.g., dlogz < 0.1; Higson et al., 2018; Speagle, 2020).
@@ II.4 Dingo-IS
The Dingo network architecture is the one presented in Dax et al. (2021), with hyperparameters kept unchanged. In short, Dingo consists of two neural networks: an embedding network and a normalizing flow. The embedding network compresses the input strain data to a set of 128 features, while the normalizing flows generate the Bayesian posterior conditioned on these features. Together, the two networks have more than $3\times 10^{8}$ learnable parameters.
To train Dingo, we generate 5 million waveforms with parameters extracted from the priors listed in Tab. 1 and the detector setup described in Sec. II.1. To speed up training and inference, we simplify the process by avoiding the iterative group-equivariant NPE (GNPE) algorithm (Dax et al., 2021, see also Sec. IV.2) and fine-tuning for varying noise ASDs, since it is fixed in our case. We also analytically sample the phase parameter $\phi_{c}$ after training rather than trying to learn it. This is similar to phase marginalization (Veitch et al., 2015; Thrane and Talbot, 2019) but does not require any approximations (Dax et al., 2023).
During training, we set aside 2% of the dataset for testing. To reduce overfitting, we use a scheduler that lowers the learning rate when the loss function stops improving after ten epochs (Paszke et al., 2019). As shown in Fig. 3, the training and test losses are closely matched after $\sim 200$ epochs, indicating no overfitting.
After training, Dingo can rapidly generate approximate posterior samples for any observed data. However, the results may deviate from the true posterior due to epistemic uncertainty, which arises when the neural network is either undertrained or not flexible enough (Kendall and Gal, 2017).
A practical way to check and refine results is through importance sampling (Tokdar and Kass, 2010; Elvira and Martino, 2021). Combining this method with Dingo leads to Dingo-IS (Dax et al., 2023). It works by assigning weights to a set of $n$ samples $\theta_{i}$ drawn from the distribution $q(\theta|d)$. The weight for each sample is calculated as follows:
@@
where $\mathcal{L}(d|\theta_{i})$ is the likelihood function, as defined in Eq. 9. In an ideal case, all the weights $w_{i}$ would be equal. However, in practice, the number of effective samples depends on the variance of the weights (Kong, 1992), and it is given by:
@@
The Bayesian evidence $\mathcal{Z}(d)$ in Eq. 1 and its uncertainty can be estimated by averaging the weights, as follows (Owen, 2013):
@@
The evidence is not only useful for model comparison, as in the standard Bayesian analysis, but can also be used to further validate NPE results. For example, the evidence would be biased if the support of the target posterior were not fully covered by $q(\theta|d)$ (Owen, 2013; Dax et al., 2023).
@@ II.5 GWFish+Priors
The FIM formalism is commonly used to simulate the PE performance of 3G detectors (Cutler and Flanagan, 1994; Vallisneri, 2008; Chan et al., 2018; Grimm and Harms, 2020; Borhanian, 2021; Iacovelli et al., 2022; Branchesi et al., 2023). In this framework, the likelihood in Eq. 1 is approximated as a multivariate Gaussian distribution centered at given injected parameters $\theta^{\mathrm{inj}}$:
@@
$\mathcal{F}_{ij}=\left(\frac{\partial h(f,\theta)}{\partial\theta_{i}}\Bigg{|}\frac{\partial h(f,\theta)}{\partial\theta_{j}}\right)\Bigg{|}_{\theta=\theta^{\mathrm{inj}}},$ (16)
with the noise-weighted inner product (|)conditional(...|...)( … | … ) defined as in Eq. 10.
We use the implementation of the FIM formalism provided in GWFish+Priors (Dupletsa et al., 2025), where posterior samples randomly drawn from Eq. 15 are weighted proportionally to the chosen prior distribution in Tab. 1.
The noise-weighted inner product (|)conditional(...|...)( … | … ), defined in Eq. 10, is used to compute the optimal SNR:
$\rho=\sqrt{(h(f,\theta^{\mathrm{inj}})|h(f,\theta^{\mathrm{inj}}))}.$ (17)
@@ III.1 Single event
Parameter $\theta^{\mathrm{inj}}$ JSD×$\times 10^{3}$ $\theta$
$\mathcal{M}_{\mathrm{d}}$ 503.7 M 1.0 nat ${516}_{-19}^{+16}$ M
$q$ 0.43 1.3 nat ${0.47}_{-0.04}^{+0.04}$
$d_{\mathrm{L}}$ 20499.4 Mpc 1.2 nat ${20000}_{-3000}^{+2000}$ Mpc
ra 2.12 rad 0.7 nat ${2.43}_{-1.82}^{+3.10}$ rad
dec -1.17 rad 1.1 nat ${-0.03}_{-1.16}^{+1.22}$ rad
$\theta_{\mathrm{jn}}$ 2.21 rad 0.6 nat ${1.28}_{-0.33}^{+0.91}$ rad
$\phi_{c}$ 3.36 rad 0.5 nat ${3.76}_{-0.33}^{+0.26}$ rad
$\psi$ 2.18 rad 1.2 nat ${1.59}_{-0.98}^{+0.95}$ rad
$t_{\mathrm{geocent}}$ 0.00005 s 1.4 nat ${0.015}_{-0.011}^{+0.012}$ s
$\chi_{1}$ 0.38 0.6 nat ${0.39}_{-0.09}^{+0.09}$
$\chi_{2}$0.39 0.6 nat ${-0.12}_{-0.29}^{+0.18}$
@@
Table 2: Injected parameters $\theta^{\mathrm{inj}}$ for Event 1 and Jensen-Shannon divergence (JSD) quantifying the deviation between Dingo-IS and Bilby for one-dimensional marginal posteriors. The last column on the right shows the median and 90% credible interval of the recovered parameters $\theta$ with Dingo-IS. We report in the bottom rows the source-frame primary and secondary mass, redshift and optimal signal-to-noise ratio ($\rho$, see Eq. 17). See Sec. III.1 for details.
@@
Figure 5: Marginalized one- and two-dimensional posterior distributions for Event 1 over all set of parameters, comparing Dingo-IS (light blue), Bilby (orange), and GWFish+Priors (black). Vertical and horizontal lines mark the true injected values (see Tab. 2). Contours represent 68% and 95% credible regions. Sky modes are discussed in Sec. IV.1. See Sec. III.1 for other details.
@@
We first validate Dingo-IS, trained on ET-$\Delta$, by comparing its results with those obtained using the standard inference code Bilby for a randomly selected event from the prior distribution, referred to hereafter as Event 1. The injected and recovered parameters for this event are listed in Tab. 2, and the corresponding signal and noise realization are shown in Fig. 2. A second event, Event 2, with a higher injected redshift and lower optimal SNR, is analyzed in Appendix B, leading to similar conclusions.
@@
For this injection, we draw $3\times 10^{4}$ samples from $q(\theta|d)$. We then analytically sample the phase $\phi_{c}$ and evaluate the importance weights (Eq. 11), which are shown in Fig. 4. The sample efficiency of this event is equal to $\sim 16\%$, corresponding to $5\times 10^{3}$ effective samples.
@@
From the importance weights shown in Fig. 4, we observe a tail with high $q(\theta|d)$ and low $p(\theta|d)$, without a corresponding tail with high $p(\theta|d)$ and low $q(\theta|d)$. This suggests that the Dingo proposal $q(\theta|d)$ covers the full support of the posterior distribution and slightly overestimates its width, which is a desired characteristic of NPE. This is because an overly broad posterior approximation results in a much smaller loss of sample efficiency compared to one that is too narrow (Dax et al., 2023).
@@
In Bilby we set the number of live points to nlive = 5000. The log evidence calculated using Dingo-IS and Bilby is $\log\mathcal{Z}(d)=-6073.89\pm 0.01$ and $-6073.58\pm 0.08$, respectively. This small difference is within statistical uncertainty.
@@
We compare posterior samples from Dingo-IS and Bilby using the Jensen–Shannon divergence (JSD; Lin, 1991), which is a symmetric extension of the Kullback–Leibler divergence (Kullback and Leibler, 1951). JSD quantifies the difference between two probability distributions, on a scale from 0 nat (identical distributions) to ln(2) = 0.69 nat (maximally distinct distributions), where the nat unit arises from using the natural logarithm. For one-dimensional posteriors, a value above $2\times 10^{-3}$ nat is considered statistically significant (Romero-Shaw et al., 2020). As shown in Tab. 2, the evaluated JSDs over marginal one-dimensional posteriors indicate that the two distributions are effectively identical. This perfect agreement is further illustrated in Fig. 5, which presents the marginalized one- and two-dimensional posteriors for all parameters.
@@
Fig. 5 shows the multimodalities and degeneracies that arise due to the geometry of a triangular co-located detector (Singh and Bulik, 2021, 2022). These effects appear in parameters such as the right ascension (ra), declination (dec), inclination angle ($\theta_{\mathrm{jn}}$), polarization angle ($\psi$) and geocentric time ($t_{\mathrm{geocent}}$). In particular, Sec. IV.1 discusses the degeneracies in sky localization, which lead to eight distinct sky modes in the two-dimensional marginal distribution of ra and dec. These modes are absent by construction when using the FIM approximation for PE.

### v2

@@ II.2 Priors
$\cos$ dec $\mathcal{U}(-1,1)$
$\sin\theta_{\mathrm{jn}}$ $\mathcal{U}(0,1)$
@@
The prior distributions $\pi(\theta)$ adopted in Eq. 1 are reported in Table 1. In particular, for the detector-frame chirp mass and mass ratio, we choose a prior uniform in primary and secondary masses (GWFish_Priors_Tutorial), therefore:
@@
where $d_{\mathrm{L}}\in[5\times 10^{3},~5\times 10^{5}]$ Mpc. For the aligned spin component $\chi_{i}$, we form a joint prior as follows:
@@
where $\pi(a_{i})=\mathcal{U}(0,0.9)$ and $\cos\theta_{i}=\mathcal{U}(-1,1)$ (lange2018rapidaccurateparameterinference; Ashton:2018jfp).
By convention, we set the Earth orientation and the positions and orientations of the three interferometers to match those at the reference time $t_{\mathrm{ref}}=1126259462.391$ s, which corresponds to the GPS trigger time of GW150914 (Romero-Shaw:2020owr). The merger time ($t_{\mathrm{geocent}}$) of all events is sampled randomly around this reference time. Fixing $t_{\mathrm{ref}}$ for such short signals is not a loss of generality, since any difference in time can be recovered.
@@ II.3 Bilby
We compare the results obtained with NPE against Bilby (Romero-Shaw:2020owr), which is a Python library specifically designed for Bayesian inference. Assuming the noise is stationary and Gaussian, the likelihood in Eq. 1 follows (Finn:1992wt):
$\mathcal{L}(d|\theta)\propto\exp\left(-\frac{1}{2}(d-h(\theta)|d-h(\theta))\right),$ (9)
where $h(\theta)$ represents the gravitational wave signal as a function of frequency $f$ for a given set of parameters $\theta$, projected onto the antenna pattern of the detector (see Sec. IV.1 for details). Eq. 9 involves the noise-weighted inner product:
$(a|b)=4~\mathrm{Re}~\int_{f_{\mathrm{min}}}^{f_{\mathrm{max}}}\mathrm{d}f\frac{a^{*}(f)b(f)}{S(f)},$ (10)
where $\mathrm{Re}$ denotes real part and complex conjugate. Moreover, $\sqrt{S(f)}$ is the noise ASD, as described in Sec. II.1, which takes into account that GW detectors are not equally sensitive at all frequencies.
To sample from Eq. 1 with the likelihood defined in Eq. 9, we use the nested sampling algorithm nessai (nessai; 2021PhRvD.103j3006W; 2023MLS&T...4c5011W), which offers fast convergence by being accelerated with normalizing flows.
Nested sampling methods produce posterior samples as a result of calculating the evidence $\mathcal{Z}(d)$ (Skilling:2004pqw; Skilling:2006gxv; sivia2006data; Buchner:2021kpm; Ashton:2022grj). Initially, a set of live points is drawn from the prior distribution. During each iteration, the live point with the lowest likelihood is replaced by a new sample. The algorithm ends once the remaining prior volume contributes less than a predefined fraction of the evidence (e.g., dlogz < 0.1; Higson_2018; Speagle:2019ivv).
@@ II.4 Dingo-IS
The Dingo network architecture is the one presented in 2021PhRvL.127x1103D, with hyperparameters kept unchanged. In short, Dingo consists of two neural networks: an embedding network and a normalizing flow. The embedding network compresses the input strain data to a set of 128 features, while the normalizing flows generate the Bayesian posterior conditioned on these features. Together, the two networks have more than $3\times 10^{8}$ learnable parameters.
To train Dingo, we generate 5 million waveforms with parameters extracted from the priors listed in Table 1 and the detector setup described in Sec. II.1. To speed up training and inference, we simplify the process by avoiding the iterative group-equivariant NPE (GNPE) algorithm (Dax:2021myb, see also Sec. IV.3) and fine-tuning for varying noise ASDs, since it is fixed in our case. We also analytically sample the phase parameter $\phi_{c}$ after training rather than trying to learn it. This is similar to phase marginalization (Veitch:2014wba; Thrane:2018qnx) but does not require any approximations (2023PhRvL.130q1403D).
During training, we set aside 2% of the dataset for testing. To reduce overfitting, we use a scheduler that lowers the learning rate when the loss function stops improving after ten epochs (NEURIPS2019_9015). As shown in Fig. 3, the training and test losses are closely matched after $\sim 200$ epochs, indicating no overfitting.
After training, Dingo can rapidly generate approximate posterior samples for any observed data. However, the results may deviate from the true posterior due to epistemic uncertainty, which arises when the neural network is either undertrained or not flexible enough (Kendall:2017tnb).
A practical way to check and refine results is through importance sampling (IS, https://doi.org/10.1002/wics.56; Elvira_2021). Combining this method with Dingo leads to Dingo-IS (2023PhRvL.130q1403D). It works by assigning weights to a set of $n$ samples $\theta_{i}$ drawn from the distribution $q(\theta|d)$. The weight for each sample is calculated as follows:
@@
where $\mathcal{L}(d|\theta_{i})$ is the likelihood function, as defined in Eq. 9. In an ideal case, all the weights $w_{i}$ would be equal. However, in practice, the number of effective samples depends on the variance of the weights (Kong), and it is given by:
@@
The Bayesian evidence $\mathcal{Z}(d)$ in Eq. 1 and its uncertainty can be estimated by averaging the weights, as follows (mcbook):
@@
The evidence is not only useful for model comparison, as in the standard Bayesian analysis, but can also be used to further validate NPE results. For example, the evidence would be biased if the support of the target posterior were not fully covered by $q(\theta|d)$ (mcbook; 2023PhRvL.130q1403D).
@@ II.5 GWFish+Priors
The FIM formalism is commonly used to simulate the PE performance of 3G detectors (Cutler:1994ys; 2008PhRvD..77d2001V; Chan:2018csa; Grimm:2020ivq; Borhanian:2020ypi; Iacovelli:2022bbs; Branchesi:2023mws). In this framework, the likelihood in Eq. 1 is approximated as a multivariate Gaussian distribution centered at given injected parameters $\theta^{\mathrm{inj}}$:
@@
$\mathcal{F}_{ij}=\left(\frac{\partial h(\theta)}{\partial\theta_{i}}\Bigg|\frac{\partial h(\theta)}{\partial\theta_{j}}\right)\Bigg|_{\theta=\theta^{\mathrm{inj}}},$ (16)
with the noise-weighted inner product (|)(...|...) defined as in Eq. 10.
We use the implementation of the FIM formalism provided in GWFish+Priors (Dupletsa:2024gfl), where posterior samples randomly drawn from Eq. 15 are weighted proportionally to the chosen prior distribution in Table 1.
The noise-weighted inner product (|)(...|...), defined in Eq. 10, is used to compute the optimal SNR:
$\rho=\sqrt{(h(\theta^{\mathrm{inj}})|h(\theta^{\mathrm{inj}}))}.$ (17)
@@ III.1 Single event
Parameter $\theta^{\mathrm{inj}}$ $\langle\mathrm{JSD}\rangle$ JSDthr $\theta$
$\mathcal{M}_{\mathrm{d}}$ 503.7 M 0.6 1.7 ${516}_{-19}^{+16}$ M
$q$ 0.43 0.9 2.1 ${0.47}_{-0.04}^{+0.04}$
$d_{\mathrm{L}}$ 20499.4 Mpc 0.9 2.0 ${20000}_{-3000}^{+2000}$ Mpc
ra 2.12 rad 0.7 1.7 ${2.43}_{-1.82}^{+3.10}$ rad
dec -1.17 rad 0.7 1.5 ${-0.03}_{-1.16}^{+1.22}$ rad
$\theta_{\mathrm{jn}}$ 2.21 rad 0.7 1.4 ${1.28}_{-0.33}^{+0.91}$ rad
$\phi_{c}$ 3.36 rad 0.4 1.4 ${3.76}_{-0.33}^{+0.26}$ rad
$\psi$ 2.18 rad 1.1 2.5 ${1.59}_{-0.98}^{+0.95}$ rad
$t_{\mathrm{geocent}}$ 0.00005 s 0.1 0.5 ${0.015}_{-0.011}^{+0.012}$ s
$\chi_{1}$ 0.38 0.3 1.2 ${0.39}_{-0.09}^{+0.09}$
$\chi_{2}$ -0.39 0.9 2.7 ${-0.12}_{-0.29}^{+0.18}$
@@
Table 2: Injected parameters $\theta^{\mathrm{inj}}$ for Event 1, along with the median Jensen-Shannon divergence ($\langle\mathrm{JSD}\rangle$, in units of $10^{-3}$ nat), which quantifies the deviation between Dingo-IS and Bilby for one-dimensional marginal posteriors. For comparison, the corresponding JSD threshold (JSDthr) is also shown. The last column on the right shows the median and 90% credible interval of the recovered parameters $\theta$ with Dingo-IS. We report in the bottom rows the source-frame primary and secondary mass, redshift, the optimal signal-to-noise ratio ($\rho$, see Eq. 17), and the sky localization error. See Sec. III.1 for details.
@@
Figure 5: Marginalized one- and two-dimensional posterior distributions for Event 1 over all sets of parameters, comparing Dingo-IS (light blue), Bilby (orange), and GWFish+Priors (black). Vertical and horizontal lines mark the true injected values (see Table 2). Contours represent 68% and 95% credible regions. Sky modes are discussed in Sec. IV.1. See Sec. III.1 for other details.
@@
We first validate Dingo-IS, trained on ET-$\Delta$, by comparing its results with those obtained using the standard inference code Bilby for a randomly selected event from the prior distribution, referred to hereafter as Event 1. The injected and recovered parameters for this event are listed in Table 2, and the corresponding signal and noise realizations are shown in Fig. 2. A second event, Event 2, with a higher injected redshift and lower optimal SNR, is analyzed in Appendix B, leading to similar conclusions.
@@
For this injection, we draw $3\times 10^{4}$ samples from the proposal distribution $q(\theta|d)$. We then analytically sample the phase $\phi_{c}$ and evaluate the importance weights (Eq. 11), which are shown in Fig. 4. The sample efficiency of this event is equal to $\sim 16\%$, corresponding to $5\times 10^{3}$ effective samples.
@@
From the importance weights shown in Fig. 4, we observe a tail with high $q(\theta|d)$ and low $p(\theta|d)$, without a corresponding tail with high $p(\theta|d)$ and low $q(\theta|d)$. This suggests that the Dingo proposal $q(\theta|d)$ covers the full support of the posterior distribution and slightly overestimates its width, which is a desired characteristic of NPE. This is because an overly broad posterior approximation results in a much smaller loss of sample efficiency compared to one that is too narrow (2023PhRvL.130q1403D).
@@
In Bilby we set the number of live points to nlive = 5000. The log evidence calculated using Dingo-IS and Bilby is $\log\mathcal{Z}(d)=-6073.89\pm 0.01$ and $6073.58\pm 0.08$, respectively. This small difference is within statistical uncertainty.
@@
We compare posterior samples from Dingo-IS and Bilby using the Jensen–Shannon divergence (JSD; 61115), which is a symmetric extension of the Kullback–Leibler divergence (10.1214/aoms/1177729694). JSD quantifies the difference between two probability distributions, on a scale from 0 nat (identical distributions) to ln(2) = 0.69 nat (maximally distinct distributions), where the nat unit arises from using the natural logarithm.
To establish a threshold for the JSD above which one-dimensional posteriors are considered to be statistically different, we proceed as follows. We use Dingo-IS to generate 100 sets of $10^{4}$ posterior samples. For each parameter $\theta$, we calculate the JSD for all 4950 unique pairs of these sets. Since all these samples are drawn from the same distribution, the resulting JSD values reflect statistical fluctuations rather than meaningful differences. We then take the 95th percentile of the resulting JSD values as a threshold for each parameter (JSDthr in Table 2). Finally, we compare the one-dimensional posteriors obtained with Bilby to each of the Dingo-IS sets, and we check whether the median of these JSD values is below the corresponding threshold ($\langle\mathrm{JSD}\rangle$ in Table 2).
As shown in Table 2, the evaluated JSDs over marginal one-dimensional posteriors indicate that the two distributions are effectively identical. This perfect agreement is further illustrated in Fig. 5, which presents the marginalized one- and two-dimensional posteriors for all parameters.
@@
Figure 5 shows the multimodalities and degeneracies that arise due to the geometry of a triangular colocated detector (Singh:2020wsy; Singh:2021bwn). These effects appear in parameters such as the right ascension (ra), declination (dec), inclination angle ($\theta_{\mathrm{jn}}$), polarization angle ($\psi$) and geocentric time ($t_{\mathrm{geocent}}$, see also Sec. IV.2). In particular, Sec. IV.1 discusses the degeneracies in sky localization, which lead to eight distinct sky modes in the two-dimensional marginal distribution of ra and dec. These modes are absent by construction when using the FIM approximation for PE.